OpenStudy (anonymous):

Find the vertex and parabola by writing in vertex form. F(x)=x^2-2x-10

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

derivative?

OpenStudy (anonymous):

first coordinate of the vertex is \(-\frac{b}{2a}\) which in your case is \(-\frac{-2}{2}=1\)

OpenStudy (anonymous):

second coordinate of the vertex is what you get when you replace \(x\) by \(1\)

OpenStudy (anonymous):

to write it in "vertex form" put \[x^2-2x-10=(x-1)^2+k\] where \(k\) is the second coordinate of the vertex found above

OpenStudy (anonymous):

like heck using calculus to find the vertex of a quadratic is like using a shotgun on a fly plus, my guess is whoever asked the question isn't in calculus, because it asks for "vertex form"

OpenStudy (anonymous):

in any case if you take the derivative of \(ax^2+bx+c\) you get \(2ax+b\) which, if you set equal to zero and solve, gives \(x=-\frac{b}{2a}\) it never changes, so you might as well just go with \(-\frac{b}{2a}\)

OpenStudy (tkhunny):

@Loser66 No. Some people are just polite and always say "Thank you". There is no natural pathway from the derivative to the vertex form. This is very clearly NOT a calculus question. The vertex form is a clear predecessor to any calculus application.

OpenStudy (loser66):

@tkhunny , I just give out the way to find vertex, I didn't indicate about the form. Yes, My fault. I apologied.

OpenStudy (tkhunny):

No worries. Always grateful for those willing to help.